OpenStudy (anonymous):

Order the Operations: Name the operation you should perform first. 1. 5+6 times 2=? 2. 18 divided 3 -1=? 3.4+(7-1)=? 4.6+6=? 5.(15+38) times 6+=? 6.5 times 10 -12=?

OpenStudy (anonymous):

use PEMDAS

OpenStudy (anonymous):

Parenthesis first.. then exponents.. then multiply.. then divite.. then add.. then subtract

OpenStudy (anonymous):

divide. not whatever i wrote

OpenStudy (anonymous):

Thank you :D

OpenStudy (anonymous):

so for #1 5+6(2) = 5 + 12 = 17

OpenStudy (anonymous):

np :) please medal if i helped u